• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Lindgren

    Log entry time 16:10:36 on February19,2007

    Entry number 194259

    keyword=Shift Summary

    Day Shift Begins
    Shift Workers
    R. Lindgren SL
    Huan Yao TO
    Runs 6592/26578 start 8:04 2.0 ua Li target 43 deg stop 8:27 1M events

    Changed target to BC
    Decrease beam to 1 ua
    Beam off for about 30 minutes due to machine problem
    10:23 Changed target to carbon
    Increase beam to 8 ua
    Dead time a little high on LHRS 13% eventhough rate is 1KHz
    Lower beam current to 6 ua deadtime dropped to 9%
    New day shift run plan in order
    see white board in counting house
    14:09 hanging angle to 78 degrees
    14:09 Started taking Cosmics
    changed angle to 78 degrees
    14:30 Stop taking Cosmics
    Start calorimeter run at .5 ua

    Runs 6609/26595 Calorimeter runs at 0.5 ua
    BcmLog_122


    Runs 6592/26578 start 8:04 2.0 ua Li target 43 deg stop 8:27 1M events
    Runs 6593/26579 start 8:30 2.0 ua Li target 43 deg stop 8:54 1M events
    Runs 6594/26580 start 9:14 1.0 ua BC target 43 deg stop 9:44 1M events
    Runs 6595/26581 start 9:43 1.0 ua BC target 43 deg stop 10:22 1M event
    Runs 6599/26585 start 10:42 6.0 ua C target 43 deg stop 11:08 1M event
    Runs 6600/26586 start 11:08 6.0 ua C target 43 deg stop 11:35 1M event
    Runs 6602/26588 start 11:42 10. ua C Target 43 deg stop 12:07 .9M
    Runs 6603/26589 start 12:08 1 ua C target 43 deg stop 12:30 .1M
    Runs 6604/26590 start 12:32 cosmics ------------- stop 12:55 13,852
    Runs 6605/26591 start 12:55 1 ua Ta target 43 deg stop 13:20 23,000
    Runs 6606/25592 start 13:22 10 ua Ta target 43 deg stop 13:44 105,000
    Runs 6607/25593 start 13:46 20 ua Ta target 43 deg stop 14:08 200,000
    Runs 6608/25594 start 14:09 cosmics ------------- stop 14:30 12,000
    Runs 6609/26595 start 14:46 .5 ua Calorimeter calib stop 14:59 87,000
    Runs 6610/26596 start 14:46 4.5 ua BC target 78 deg stop 16:07 0.2M

    Day Shift ends